What is the discount yield, bond equivalent yield, and effective annual return on a $1 million commercial paper issue that currently sells at 96.375 percent of its face value and is 70 days from maturity? (Use 360 days for discount yield and 365 days in a year for bond equivalent yield and effective annual return. Do not round intermediate calculations. Round your answers to 3 decimal places. (e.g., 32.161))

1.) Discount Yield % (Round all answers to 3 decimal places)

2.) Bond equivalent Yield %

3.) Effective annual return %
Answer #1

1.) Discount Yield = [($1,000,000 − $963,750) / $1,000,000](360/70) = 18.643%

2.) Bond equivalent Yield = [($1,000,000 − $963,750) / $963,750](365/70) = 19.613%

3.) Effective annual return = (1 + 0.19613/(365/70))365/70 − 1 = 21.231%